OpenStudy (anonymous):

10 1/6+ 5 1/2= 15 2/3?

OpenStudy (mathteacher1729):

Ok, so before adding mixed numbers -- do you know how add fractions? For example -- can you add 1/6 + 1/2 ?

OpenStudy (anonymous):

wait its wrong? and yes 4/6

OpenStudy (mathteacher1729):

How did you get 4/6 ? Did you use a calc or do it by hand, etc?

OpenStudy (mathteacher1729):

(1/2 + 1/6 = 4/6 is correct, by the way, I just wanna know how you're approaching the problem.)

OpenStudy (anonymous):

by hand: 1/6 is can stay as it is because 1/2 can be multiplied by 3: Now you have 1/6+ 3/6= 4/6

OpenStudy (anonymous):

wouldn't you do; 10+5=15 next, and reduce 4/6 to 2/3???

OpenStudy (mathteacher1729):

Yup. Exactly. :) Now to convert a number like 10 into a fraction you can write it as 10/1 or 20/2 or 30/3 etc... all of those equal 10.

OpenStudy (anonymous):

Thanks!(;

OpenStudy (mathteacher1729):

So 10 1/6 is really saying " 10 plus 1/6". and 5 1/2 is really saying "5 plus 1/2" and 15 2/3 is really saying " 15 plus 2/3" Can you convert each of those into fractions?

OpenStudy (anonymous):

wouldnt be converting it int improper fractions example 10x6+! or 5x2+1 and 15x3+2